OpenStudy (anonymous):

Solve for y x=(1-2y)/(3y) I can't seem to figure out how to get those y's over to the other side :(

OpenStudy (anonymous):

first, multiply both sides by 3y, which gives you 3xy = 1 - 2y then add 2y to both sides, which gives you 3xy + 2y = 1 then factor out the y from both terms on the left side of the equal sign y(3x + 2) = 1 last step, divide both sides by (3x + 2), which gives you y = 1/(3x + 2)